# Generating Plausible Distractors for Multiple-Choice Questions via Student Choice Prediction

### v2

@@ Abstract
In designing multiple-choice questions (MCQs) in education, creating plausible distractors is crucial for identifying students’ misconceptions and gaps in knowledge and accurately assessing their understanding. However, prior studies on distractor generation have not paid sufficient attention to enhancing the difficulty of distractors, resulting in reduced effectiveness of MCQs. This study presents a pipeline for training a model to generate distractors that are more likely to be selected by students. First, we train a pairwise ranker to reason about students’ misconceptions and assess the relative plausibility of two distractors. Using this model, we create a dataset of pairwise distractor ranks and then train a distractor generator via Direct Preference Optimization (DPO) to generate more plausible distractors. Experiments on computer science subjects (Python, DB, MLDL) demonstrate that our pairwise ranker effectively identifies students’ potential misunderstandings and achieves ranking accuracy comparable to human experts. Furthermore, our distractor generator outperforms several baselines in generating plausible distractors and produces questions with a higher item discrimination index (DI). [^1] This paper is currently under review. All source code, fine-tuned models, and a subset of our data will be made publicly available upon publication.
@@
Yooseop Lee1  Suin Kim2  Yohan Jo1thanks: Corresponding author. 1Seoul National University 2Elice {lyooseop, yohan.jo}@snu.ac.krsuin@elicer.com
@@ 1 Introduction
Based on these needs, this study presents a model training pipeline for distractor generation. Figure 1 illustrates example distractors generated by GPT and our model. Our main idea is to assign relative ranks to distractors based on which ones students are more likely to select, and use this information to train a model to generate plausible distractors. To achieve this, the process involves three steps (Figure 2). First, we train a pairwise ranker to predict which distractors are more plausible and likely to confuse students (Step 1). Next, we create a synthetic student choice dataset that includes pairwise ranking information among distractors (Step 2). Finally, leveraging this dataset, we train a distractor generator by applying Direct Preference Optimization (DPO, Rafailov et al., 2024) (Step 3).
@@ 2.2 Pairwise Ranker
Our study aims to assign plausibility ranks among distractors using an LLM (Figure 2, Step 1 and 2). This approach is motivated by prior findings demonstrating that LLMs exhibit strong inferential abilities, closely aligning with human performance in many evaluation tasks (Sun et al., 2023; Liu et al., 2023). Moreover, distilling these abilities from LLMs into smaller models, such as Prometheus 2 (Kim et al., 2024), fine-tuned from Mistral (Jiang et al., 2023), and ListT5 (Yoon et al., 2024), has achieved comparable performance to LLMs while offering faster inference and reduced positional biases.
However, the reasoning abilities of LLMs to rank plausible distractors remain underexplored. A related study by Scarlatos et al. (2024) proposed an approach that trains a pairwise ranker for distractors using data on the actual selection rates of distractors by students. They further applied DPO to prioritize more plausible distractors. However, their model neither examines nor leverages LLMs’ reasoning abilities, and the trained model lacks interpretability. In contrast, our study extensively evaluates LLMs’ reasoning abilities by comparing various prompting approaches that are broadly applicable across diverse subjects. Additionally, our ranker generates reasoning behind its choices, enhancing its interpretability.
@@ 3.1 Base MCQ Dataset
A key feature of this dataset is that it includes information on how many students answered each question and the selection rate for each distractor. This allows us to determine which distractors were more confusing and plausible to students. Since each question was solved by hundreds of students from diverse sectors, the selection rate information is considered reliable. This information will play a key role in training the pairwise ranker and distractor generator, as discussed later. We will release a subset of this dataset—52 questions with no licensing issues—to the community.
@@ 5.2 Distractor Generator
For English questions, we used 88 questions from a South Korean high school exam dataset to train a pairwise ranker and a distractor generator. In Table 7, our pairwise ranker, despite limited training data, outperformed GPT-3.5-turbo and closely approached GPT-4o. Similarly, Table 6 shows that in Setting B, where more distractors were compared, our DPO model achieved higher plausibility than GPT models, reflecting the trends observed in CS subjects.
@@ A.2 SFT and DPO Settings for Pairwise Ranker
The pairwise ranker model was trained using Mistral-7B-Instruct-v0.2333This model is distributed under the Apache 2.0 license. https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 with 4-bit quantization and fine-tuned using LoRA. For SFT, the learning rate was set to 2e-4 and the model was trained for 5 epochs. For DPO, the learning rate was set to 1e-6, also trained for 5 epochs. These hyperparameters were selected as they allowed stable training without overfitting while preserving the quality of the DPO output. SFT took approximately 2 hours, and DPO took about 1 hour on an NVIDIA A6000 GPU. Scarlatos et al. (2024) model was reproduced for baseline comparison using the same model and DPO settings as above.

### v3

@@ Abstract
In designing multiple-choice questions (MCQs) in education, creating plausible distractors is crucial for identifying students’ misconceptions and gaps in knowledge and accurately assessing their understanding. However, prior studies on distractor generation have not paid sufficient attention to enhancing the difficulty of distractors, resulting in reduced effectiveness of MCQs. This study presents a pipeline for training a model to generate distractors that are more likely to be selected by students. First, we train a pairwise ranker to reason about students’ misconceptions and assess the relative plausibility of two distractors. Using this model, we create a dataset of pairwise distractor ranks and then train a distractor generator via Direct Preference Optimization (DPO) to generate more plausible distractors. Experiments on computer science subjects (Python, DB, MLDL) demonstrate that our pairwise ranker effectively identifies students’ potential misunderstandings and achieves ranking accuracy comparable to human experts. Furthermore, our distractor generator outperforms several baselines in generating plausible distractors and produces questions with a higher item discrimination index (DI).  [^1] Our code and a subset of our dataset are available at https://github.com/holi-lab/distractor-generator
@@
Yooseop Lee1,2thanks: This work was conducted while the first author was a graduate student at Seoul National University.  Suin Kim2  Yohan Jo1thanks: Corresponding author. 1Graduate School of Data Science, Seoul National University 2Elice {lyooseop, yohan.jo}@snu.ac.krsuin@elicer.com
@@ 1 Introduction
Based on these needs, this study presents a model training pipeline for distractor generation. Figure 1 illustrates example distractors generated by GPT and our model. Our main idea is to assign relative ranks to distractors according to their likelihood of being selected by students, and use this information to train a model to generate plausible distractors. To achieve this, the process involves three steps (Figure 2). First, we train a pairwise ranker to predict which distractors are more plausible and likely to confuse students (Step 1). Next, we create a synthetic student choice dataset that includes pairwise ranking information among distractors (Step 2). Finally, leveraging this dataset, we train a distractor generator by applying Direct Preference Optimization (DPO, Rafailov et al., 2024) (Step 3).
@@ 2.2 Pairwise Ranker
Our study aims to assign plausibility ranks among distractors using an LLM (Figure 2, Step 1 and 2). This approach is motivated by prior findings demonstrating that LLMs exhibit strong inferential abilities, closely aligning with human performance in many evaluation tasks (Sun et al., 2023; Liu et al., 2023). Moreover, distilling these abilities from LLMs into smaller models, such as Prometheus 2 (Kim et al., 2024) fine-tuned from Mistral (Jiang et al., 2023), and ListT5 (Yoon et al., 2024), has achieved comparable performance to LLMs while offering faster inference and reduced positional biases.
However, the reasoning abilities of LLMs to rank plausible distractors in the education domain remain underexplored. A related study by Scarlatos et al. (2024) proposed an approach that trains a pairwise ranker for distractors using data on the actual selection rates of distractors by students. They further applied DPO to prioritize more plausible distractors. However, their model neither examines nor leverages LLMs’ reasoning abilities, and the trained model lacks interpretability. In contrast, our study extensively evaluates LLMs’ reasoning abilities by comparing various prompting approaches that are broadly applicable across diverse subjects. Additionally, our ranker generates reasoning behind its choices, enhancing its interpretability.
@@ 3.1 Base MCQ Dataset
A key feature of this dataset is that it includes information on how many students answered each question and the selection rate for each distractor. This allows us to determine which distractors were more confusing and plausible to students. Since each question was solved by hundreds of students from diverse sectors, the selection rate information is considered reliable. This information will play a key role in training the pairwise ranker and distractor generator, as discussed later. A subset of this dataset without licensing issues is available on our project website (the footnotes on the first page).
@@ 5.2 Distractor Generator
For English questions, we used 88 questions from a South Korean high school exam [^3] These MCQs are from the latest CSAT (Korean SAT), and the distractor selection rates were obtained from an online education platform specializing in CSAT preparation. to train a pairwise ranker and a distractor generator. In Table 7, our pairwise ranker, despite limited training data, outperformed GPT-3.5-turbo and closely approached GPT-4o. Similarly, Table 6 shows that in Setting B, where more distractors were compared, our DPO model achieved higher plausibility than GPT models, reflecting the trends observed in CS subjects.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grants funded by the Korean government (MSIT) (RS-2024-00333484, RS-2024-00414981). It was also supported by Elice, Inc., which also provided the proprietary datasets.
@@ A.2 SFT and DPO Settings for Pairwise Ranker
The pairwise ranker model was trained using Mistral-7B-Instruct-v0.2444This model is distributed under the Apache 2.0 license. https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 with 4-bit quantization and fine-tuned using LoRA. For SFT, the learning rate was set to 2e-4 and the model was trained for 5 epochs. For DPO, the learning rate was set to 1e-6, also trained for 5 epochs. These hyperparameters were selected as they allowed stable training without overfitting while preserving the quality of the DPO output. SFT took approximately 2 hours, and DPO took about 1 hour on an NVIDIA A6000 GPU. Scarlatos et al. (2024) model was reproduced for baseline comparison using the same model and DPO settings as above.